# Two-pion exchange nucleon-nucleon potentials with Roper resonance excitation

### v1

@@ 1 Preparation
where AA and $B$ symbolize vertices on the left and right hand side of the diagram and the angular integral $\int_{-1}^{1}\!dx$ effectively goes over the Lorentz-invariant $2\pi$-phase space. Furthermore, $\mu=\sqrt{q\!\cdot\!q}\geq 2m_{\pi}$ denotes the $2\pi$-invariant mass and $k=\sqrt{\mu^{2}/4-m_{\pi}^{2}}$ is the (single) pion center-of-mass momentum. For the $\bar{N}\!N\to\pi\pi\to\bar{N}\!N$ $t$-channel kinematics under consideration in the heavy baryon limit, one sets the nucleon velocity four-vector $v^{\nu}$ equal to $v^{\nu}=(0,i\,\vec{v}\,)$ with $\vec{v}$ a 3-component unit-vector, which leads to relations $v\!\cdot\!q=0$ and $v\!\cdot\!l=-i\,x\,k$. In the first step the emerging integrand is complex-valued, but it becomes real-valued after symmetrizing with respect to $x\to-x$, and leaves an integral $\int_{0}^{1}\!dx$ over positive values of the cosine $x$. Further technical details on the method can be found in section 2.2 of ref.[1], but note that therein the numerator of eq.(9) must be corrected to $\sqrt{\mu^{2}-4m_{\pi}^{2}}$.
@@
with $g^{\prime}_{A}$ a coupling constant (various values are discussed in section III.A of ref.[2]) and $\rho\simeq 500\,$MeV the mass difference between the Roper $N^{*}(1440)$ resonance and the nucleon. The right half of eq.(4) gives the analogous input for $\Delta(1232)$-isobar excitation [3], where the $\Delta N$-mass difference is smaller: $\Delta\simeq 300\,$MeV. The $2\times 4$ spin- and isospin-transition matrices $S_{i}$ and $T_{a}$ fulfill the relations $3S_{i}S_{j}^{\dagger}=2\delta_{ij}-i\epsilon_{ijk}\sigma_{k}$ and $3T_{a}T_{b}^{\dagger}=2\delta_{ab}-i\epsilon_{abc}\tau_{c}$. The coupling ratio $3/\sqrt{2}$ to $\pi N\!N$ is empirically well satisfied.
@@ 2 Spectral functions of $2\pi$-exchange NN-potentials
As the diagram involves on one side the Weinberg-Tomozawa $2\pi$-contact vertex, there is no $g_{A}^{2}$ in the prefactor as it was incorrectly written in eq.(A1) of ref.[2].
@@
Modulo additive constants (plus const $q^{2}$) and the opposite sign-convention the results for $V_{C}(q)$ and $W_{T}(q)$ agree with the expressions written in eqs.(5,14) of ref.[2].
@@
The spectral functions for the $2\pi$-exchange NN-potentials with single and double $\Delta(1232)$-isobar excitation (computed in section 3 of ref.[3]) are obtained from the expressions written here in eqs.(5-13) by replacing the coupling constant $g^{\prime}_{A}\to g_{A}$, and multiplying with weight factors 1,2,1,1,1/2,4,1,1,1/4-1,2,-1,-1,1/2,4,1,1,1/4, in that order.
@@
As emphasized in ref.[2], when including the Roper resonance one must also consider explicit contributions from $\Delta(1232)$-isobar excitation because of the lower $\Delta N$-mass splitting. The NN-potentials as they arise from single and double $\Delta(1232)$-excitation have been given in section  3 of ref.[3]. Here, we complete the pattern of $2\pi$-exchange with intermediate low-lying nucleon-resonance excitations by considering the combination of the $\Delta(1232)$-isobar and the Roper resonance. The (right) planar and crossed box diagrams in Fig. 1 with combined $\Delta(1232)$-isobar and Roper excitation and their mirror partners lead to the following (doubled) contributions to the spectral functions:
@@
which are obviously symmetric under the permutation of the mass-splittings $\Delta\leftrightarrow\rho$. The apparent complexity of one-loop integrals with heavy baryon propagators involving different mass-splittings boils down on the level of the spectral function to symmetrizing a term involving $\arctan(k/\Delta)$ by another term involving $\arctan(k/\rho)$.
By comparison with the analytical expressions given in section 3 of ref.[3], one sees that the translation of $\mu$-dependent spectral functions into $q$-dependent $2\pi$-exchange NN-potentials proceeds as follows:

### v2

@@ 1 Preparation
where AA and $B$ symbolize vertices on the left and right hand side of the diagram and the angular integral $\int_{-1}^{1}\!dx$ effectively goes over the Lorentz-invariant $2\pi$-phase space. Furthermore, $\mu=\sqrt{q\!\cdot\!q}\geq 2m_{\pi}$ denotes the $2\pi$-invariant mass and $k=\sqrt{\mu^{2}/4-m_{\pi}^{2}}$ is the (single) pion center-of-mass momentum. For the $\bar{N}\!N\to\pi\pi\to\bar{N}\!N$ $t$-channel kinematics under consideration in the heavy baryon limit, one sets the nucleon velocity four-vector $v^{\nu}$ equal to $v^{\nu}=(0,i\,\vec{v}\,)$ with $\vec{v}$ a 3-component unit-vector, which leads to the relations $v\!\cdot\!q=0$ and $v\!\cdot\!l=-i\,x\,k$. In the first step the emerging integrand is complex-valued, but it becomes real-valued after symmetrizing it with respect to $x\to-x$, and leaves an integral $\int_{0}^{1}\!dx$ over positive values of the cosine $x$. The intermediate state nucleon is handled with an infinitesimally small positive mass-splitting $\eta$, which leads to $\lim_{\eta\to 0}\eta/(\eta^{2}+k^{2}x^{2})=\pi\delta(x)/k$ and $\int_{0}^{1}\!dx\,\delta(x)=1/2$. Further technical details on the method can be found in section 2.2 of ref.[1], but note that therein the numerator of eq.(9) must be corrected to $\sqrt{\mu^{2}-4m_{\pi}^{2}}$.
@@
with $g^{\prime}_{A}$ a coupling constant (various values are discussed in section III.A of ref.[2]) and $\rho\simeq 500\,$MeV the mass difference between the Roper $N^{*}(1440)$-resonance and the nucleon. The right half of eq.(4) gives the analogous input for $\Delta(1232)$-isobar excitation [3], where the $\Delta N$-mass difference is smaller: $\Delta\simeq 300\,$MeV. The $2\times 4$ spin- and isospin-transition matrices $S_{i}$ and $T_{a}$ fulfill the relations $3S_{i}S_{j}^{\dagger}=2\delta_{ij}-i\epsilon_{ijk}\sigma_{k}$ and $3T_{a}T_{b}^{\dagger}=2\delta_{ab}-i\epsilon_{abc}\tau_{c}$. The coupling ratio $3/\sqrt{2}$ to $\pi N\!N$ is empirically well satisfied.
@@ 2 Spectral functions of $2\pi$-exchange NN-potentials
Since this diagram involves on one side the isovector Weinberg-Tomozawa $2\pi N\!N$-contact vertex, there is no $g_{A}^{2}$ in the prefactor, as it was incorrectly written in eq.(A1) of ref.[2].
@@
Modulo additive constants (plus const $q^{2}$) and the opposite sign-convention, the results for $V_{C}(q)$ and $W_{T}(q)$ agree with the expressions written in eqs.(5,14) of ref.[2].
@@
The spectral functions for the $2\pi$-exchange NN-potentials with single and double $\Delta(1232)$-isobar excitation (computed in section 3 of ref.[3]) are obtained from the expressions written here in eqs.(5-13) by replacing the coupling constant $g^{\prime}_{A}\to g_{A}$, the mass-splitting $\rho\to\Delta$, and multiplying with weight factors 1,2,1,1,1/2,4,1,1,1/4-1,2,-1,-1,1/2,4,1,1,1/4, in that order.
@@
As emphasized in ref.[2], when including the Roper $N^{*}(1440)$-resonance one must also consider explicit contributions from $\Delta(1232)$-isobar excitation because of the lower $\Delta N$-mass splitting. The NN-potentials as they arise from single and double $\Delta(1232)$-excitation have been given in section  3 of ref.[3]. Here, we complete the pattern of $2\pi$-exchange with intermediate low-lying nucleon-resonance excitations by considering the combination of the $\Delta(1232)$-isobar and the Roper resonance. The (right) planar and crossed box diagrams in Fig. 1 with combined $\Delta(1232)$-isobar and Roper excitation and their mirror partners lead to the following (doubled) contributions to the spectral functions:
@@
which are obviously symmetric under the permutation of the mass-splittings $\Delta\leftrightarrow\rho$. The apparent complexity of one-loop integrals with heavy baryon propagators involving different mass-splittings boils down on the level of the spectral function to symmetrizing a term involving $\Delta^{-1}\!\arctan(k/\Delta)$ by another term involving $\rho^{-1}\!\arctan(k/\rho)$.
By comparison with the analytical expressions given in section 3 of ref.[3], one learns that the translation of $\mu$-dependent spectral functions into $q$-dependent $2\pi$-exchange NN-potentials proceeds as follows: